OpenStudy (anonymous):

CHECK ALL OF THE ORDERED PAIRS THAT SATISFY THE EQUATION BELOW Y=2/3 a) (14,21) b) (42,28) c) ( 21, 3/2) d) (21,14) E) (7,14) f) (21,42) ANSWER PLEASE

OpenStudy (ivettef365):

I believe you are missing the x on the equation, wouldn't it be y = 2/3x

OpenStudy (anonymous):

YEAH I AM

OpenStudy (ivettef365):

the only thing you need to do is substitute the pair (x,y) to see if is true for example first pair you have (14,21)

OpenStudy (anonymous):

DO YOU NO THE ANSWERS

OpenStudy (anonymous):

?

OpenStudy (ivettef365):

then it would be y = 2/3x 21 = 2/3 (14) 21 = 28/3 not true

OpenStudy (ivettef365):

now lets try point (21,14) y = 2/3x 14 = 2/3(21) 14 = 42/3 14 = 14 this pair is good

OpenStudy (ivettef365):

you have to check all pairs that will satisfy the equation

OpenStudy (anonymous):

YES.

OpenStudy (ivettef365):

ok, already showed you how to check, you can try the other ones

OpenStudy (anonymous):

WAIT CAN YOU TELL ME THE ANSWERS?